# Prompt-based Pseudo-labeling Strategy for Sample-Efficient Semi-Supervised Extractive Summarization

### v2

@@ 4.1 Datasets
| Method | R-2 | G-Eval | R-2 | G-Eval | R-2 | G-Eval |
| PreSumm (50 labels) | 37.1 (1.1) | 29.3 (0.6) | 3.2 (0.8) | 31.2 (1.7) | 7.3 (0.9) | 11.4 (2.3) |
| PreSumm (300 labels) | 51.1 (2.1) | 58.7 (1.3) | 7.6 (0.6) | 66.7 (1.2) | 10.8 (0.9) | 48.5 (3.4) |
| Random | 32.4 (1.9) | 31.1 (1.2) | 2.4 (1.5) | 34.7 (1.8) | 6.1 (0.2) | 14.1 (2.7) |
| Confidence + G-4 score | 42.8 (0.7) | 43.8 (0.7) | 5.6 (0.5) | 57.5 (1.4) | 7.9 (0.2) | 37.1 (1.8) |
| + G-4 relabeling (Ours) | 50.4 (0.8) | 56.3 (0.9) | 7.2 (0.6) | 69.3 (2.1) | 9.7 (1.2) | 49.3 (1.4) |
4.2 Training and Evaluation Details.
Training. We use the TransformerSum repository111https://transformersum.readthedocs.io/en/latest/ to implement our training pipeline. We use PreSumm as our teacher model and experiment with two backbones: distilbert-base-uncased and bert-base-uncased. We fix the size of the labeled set to 50 for all the datasets and set to 50. We add 5 pseudo-labels to the training set in each cycle, thus resulting in a final training set size of 300 (50 labeled + 250 pseudo-labeled examples.) We set the summary size to 4 for TweetSumm and 8 for WikiHow and ArXiv/PubMed. We base these summary sizes on the average summary size of the labeled training set.
We start with a learning rate of on all the datasets and use a cyclic learning rate scheduler during training (Smith, 2015), which is the default setting in TransformerSum. Additionally, we use AdamW as our optimizer with . We train all our models on a single V100 GPU with 12G VRAM. We repeat each experiment for three different seeds and report the mean and standard deviation in our results unless otherwise stated.
Evaluation. First, we use the standard ROUGE- (R-) and ROUGE-L (R-L) F1 scores (Lin, 2004) on the test set to evaluate summary quality. ROUGE- measures the -gram overlap between the predicted and the ground truth summaries, whereas ROUGE-L also takes into consideration the order of -grams. We use pyrouge to compute ROUGE scores in our setup and report them in Table 2.
Additionally, we use G-Eval Liu et al. (2023) for evaluation, which is an evaluation framework using an LLM like GPT-4. We use G-Eval as it substantially surpasses ROUGE scores and previous model-based evaluation metrics, such as BERTScore (Zhang et al., 2019) and BARTScore (Yuan et al., 2021), for summarization. More importantly, it is shown to be better aligned with human preferences than the existing automated evaluation metrics for text summarization. We report our G-Eval results in Table 3
4.3 Baselines
We compare our method with the following baselines: 1) PreSumm (Liu and Lapata, 2019). The original PreSumm model that pretrains a BERT model for summarization. We train two PreSumm models – one on a limited training set with 50 labeled examples to match the starting point of our semi-supervised setting and another with 300 labeled examples, the same as the dataset size at the end of our training cycle. 2) Teacher-Student Learning (TSL) (Zhuang et al., 2023). Current state-of-the-art semi-supervised method on TweetSumm. The teacher-student learning framework uses a similar formulation for computing model confidence to ours, as given below:
In Equation 3, , where is the probability of sentence being selected for summary for dialog estimated by the teacher model, and . We report the performance of the TSL (50:500) model from the paper, as it was the closest to our setup (50 labeled examples + 500 unlabeled examples). 3) Confidence + G-4 relabeling + G-4 score (Ours). Our proposed method following the methodology in Section 3. We first use the PreSumm teacher model to shortlist 50 pseudo-labels (Stage 1 and 2), relabel them using GPT-4 (Stage 3), and then select the top 5 using GPT-4 score (Stage 4). 4) Confidence + G-4 score. We skip Stage 3 from 3) to directly score the top 50 PreSumm pseudo-labels using GPT-4. We run this baseline to measure the effect of relabeling in our pipeline. 5) Confidence + G-4 relabeling. We skip Stage 4 from 3) and select the final 5 pseudo-labels based on PreSumm confidence. We run this baseline to measure the importance of the scoring mechanism. 6) Confidence. We skip Stage 3 and 4 from from 3) and select 5 PreSumm pseudo-labels based on PreSumm confidence. 7) Random. Same as 6) but instead of using teacher confidence defined in Equation 2, we randomly select five PreSumm pseudo-labels to include in each cycle.
@@ 5 Results
We now discuss results in Table 2 and Table 3. Table 2 reports the ROUGE scores for different pseudo-labeling strategies for the three datasets. All the methods operate in (50:250) setup, except TSL (50:500), which uses 500 unlabeled examples. Notably, Table 2 has two blocks of results, one for DistilBERT as the backbone of the PreSumm model and another for BERT as the PreSumm backbone. Table 3 compares G-Eval scores of our semi-supervised approach with fully-supervised setups.
@@ 5.1 Comparison of Pseudo-label Selection Strategies
Referring to Table 2, we note that all pseudo-label selection strategies outperform the random baseline. The “Random" baseline performs worse than “PreSumm (50 labels)" on all datasets (R-2 in Table 2 v/s R-2 in Table 3), meaning that the majority of the shortlisted PreSumm pseudo-labels are low-quality. Using teacher confidence leads to slight performance gains on all the datasets, and adding GPT-4 score further improves the results (“Confidence" v/s “Confidence + G-4 score" in Table 2). These improvements indicate that the shortlisted PreSumm pseudo-labels include some good-quality pseudo-labels, too, and using GPT-4 to rate those pseudo-labels is crucial to picking them.
To further confirm our findings, we conduct a qualitative study, where we compute the ROUGE scores of the final 5 pseudo-labels for each method against the respective ground truth summaries, and Figure 5 shows the mean ROUGE-2 of the five selected pseudo-labels. To clarify, we obtained the “Oracle" results by directly selecting the final 5 pseudo-labels using ROUGE-2 scores computed against the ground truth. We note a stark difference between “Confidence" and “Oracle," which shows that relying solely on teacher confidence consistently leads to a selection of low-quality pseudo-labels. Combining GPT-4 score with teacher confidence is effective (“Confidence + G-4 score"), and adding the GPT-4 relabeling greatly boosts the quality of selected pseudo-labels (“Ours").
@@ 5.2 Effect of Relabeling
Referring to Tables 2 and 3, we observe that relabeling with GPT-4 leads to a significant boost in the summarization performance in terms of both ROUGE scores and G-Eval. When using BERT as the backbone, we note that the ROUGE-1 improves from 52.3 to 58.9 on the TweetSumm dataset, 25.2 to 26.1 on the WikiHow dataset, and 27.7 to 29.1 on the ArXiv/PubMed dataset. GPT-4 relabeling is also effective when using teacher confidence without GPT-4 score (“Confidence" v/s “Confidence + G-4 relabeling"). Our previous qualitative study also supports these results, showing that relabeling improves the quality of pseudo-labels. We observe similar trends when using DistilBERT as PreSumm’s backbone.
We conduct additional testing to analyze the performance of our best- and second-best-performing models, both of which involve relabeling. We find that the p-value < 0.016 for Welch’s test for an R-1 of 58.9 (1.4) for “Confidence + G-4 score + G-4 relabelling" v/s 57.8 (1.2) for “Confidence + G-4 relabeling" on the TweetSumm dataset, denoting the differences are significant. We conduct similar tests on the other two datasets and find that the differences between best- and second-best performing models are significant.
@@ 5.3 On Sample Efficiency of Different Approaches
Referring to Table 3, we now analyze the sample efficiency of different learning algorithms. For fully supervised methods, we note that including more labeled examples improves G-Eval and ROUGE scores across the board (“PreSumm (50 labels)" v/s “PreSumm (300 labels)"). Our semi-supervised approach using 50 labels with GPT-4 relabeling and GPT-4 score achieves competitive performance to the fully supervised PreSumm model trained on 300 labels. Notably, we get better G-Eval scores than “PreSumm (300 labels)" on WikiHow and ArXiv/PubMed datasets and are competitive on TweetSumm. This is encouraging, as “PreSumm (300 labels)" approximates the best-case scenario of 100% high-quality labels in the training set.
To better understand sample efficiency, we plot the R-1 scores against the number of training cycles, as shown in Figure 6. We note that the “Random" setting is highly unstable, which further explains its worse performance than the “PreSumm (50 labels)" model. Next, “Confidence + G-4 score" slightly improves over “Confidence" on TweetSumm and WikiHow, but more importantly, it is consistently more stable. Finally, our method with GPT-4 scoring and relabeling not only significantly boosts the R-1 scores (visible gap between “Ours" and the rest) but also does so at a much faster rate. For all the datasets, our method peaks and stabilizes under 20 cycles (100 pseudo-labels), further endorsing the sample efficiency of our method compared to other approaches.
@@ 6 Conclusion
First, we rely completely on GPT’s pseudo-labels for scoring and relabelling, which is not perfect and we do not currently have a way to automatically judge them. We need a human in the loop to judge the quality of such pseudo-labels. GPT-4 is also a closed-source model, and the performance of open-source models for the task is worth exploring.
Second, we can also potentially improve the sample efficiency of the model. Currently, we relabel all the top 50 pseudo-labels selected by PreSumm in Stage 3. This is cheaper than pseudo-labeling all the pseudo-labels, but there’s a chance that GPT-4 relabels an already good-quality summary, as we note from our analysis that the initial shortlist contains some good pseudo-labels. Therefore, we can develop an additional strategy to decide whether to relabel a given pseudo-label. Such a method would further reduce costs for using closed-source models like GPT-4. We hope this work furthers research toward incorporating LLMs to address such problems for low-resource semi-supervised tasks.
@@
We use GPT to generate new examples, and even though GPT-4 is instruction-tuned, some generations might depict undesirable biases with respect to the current objective. For the stated reason, we recommend using our proposed models with human monitoring when employing the method for real-world settings. Practitioners may also consider explicitly debiasing language models for their specific use cases (Barikeri et al., 2021; Schick et al., 2021).

### v3

@@ 4.3 Baselines
Method | R-2 | L-Eval | R-2 | L-Eval | R-2 | L-Eval |
PreSumm (50 labels) | 37.1 (1.1) | 31.2 (0.5) | 3.2 (0.8) | 34.2 (1.5) | 7.3 (0.9) | 13.5 (1.2) |
PreSumm (300 labels) | 51.1 (2.1) | 60.5 (1.2) | 7.6 (0.6) | 68.1 (1.1) | 10.8 (0.9) | 49.5 (2.4) |
PreSumm (500 labels) | 54.4 (1.2) | 67.1 (0.3) | 7.9 (0.5) | 74.4 (0.6) | 11.3 (0.5) | 58.2 (1.1) |
PreSumm (750 labels) | 56.1 (0.7) | 70.3 (0.5) | 8.5 (0.4) | 76.5 (0.4) | 12.1 (0.7) | 62.8 (0.7) |
50 labels | ||||||
Random | 32.4 (1.9) | 32.1 (1.1) | 2.4 (1.5) | 37.7 (1.6) | 6.1 (0.2) | 15.1 (2.3) |
Confidence + G-4 score | 42.8 (0.7) | 46.2 (0.2) | 5.6 (0.5) | 59.4 (1.3) | 7.9 (0.2) | 40.1 (1.9) |
+ G-4 relabeling (Ours) | 50.4 (0.8) | 58.4 (0.4) | 7.2 (0.6) | 70.3 (1.4) | 9.7 (1.2) | 52.5 (1.3) |
Confidence + L-3 score | 41.6 (1.2) | 45.8 (0.7) | 5.2 (0.2) | 57.5 (1.4) | 7.9 (0.5) | 37.1 (1.8) |
+ L-3 relabeling (Ours) | 50.1 (0.3) | 56.3 (0.9) | 6.9 (0.3) | 69.3 (2.1) | 9.4 (0.7) | 49.3 (1.4) |
500 labels | ||||||
Random | 42.7 (1.1) | 52.3 (1.2) | 4.5 (1.1) | 52.7 (1.8) | 9.5 (1.2) | 44.1 (0.9) |
Confidence + L-3 score | 55.6 (0.8) | 69.2 (0.7) | 8.2 (0.5) | 75.2 (1.4) | 11.8 (0.2) | 60.2 (0.5) |
+ L-3 relabeling (Ours) | 56.2 (0.8) | 71.2 (0.9) | 8.8 (0.3) | 77.3 (1.3) | 12.3 (0.2) | 65.7 (0.3) |
@@ 5 Results
We now discuss results in Table 2 and Table 3. Table 2 reports the ROUGE scores for different pseudo-labeling strategies for the three datasets. All the methods operate in (50:250) or (500:250) setups, except TSL (50:500), which uses 500 unlabeled examples. Notably, Table 2 has two blocks of results, one for DistilBERTbase as the backbone of the PreSumm model and another for BERTbase as the PreSumm backbone. Table 3 compares L-Eval scores of our semi-supervised approach with fully-supervised setups.
@@ 5.1 Comparison of Pseudo-label Selection Strategies
Referring to Table 2, we note that all pseudo-label selection strategies outperform the random baseline. The “Random" baseline performs worse than the fully supervised counterpart on all datasets (R-2 in Table 2 v/s R-2 in Table 3), meaning that the majority of the shortlisted PreSumm pseudo-labels are low-quality. Using teacher confidence leads to slight performance gains on all the datasets, and adding GPT-4 score further improves the results (“Confidence" v/s “Confidence + G-4 score" in Table 2). These improvements indicate that the shortlisted PreSumm pseudo-labels include some good-quality pseudo-labels, too, and using GPT-4 to rate those pseudo-labels is crucial to picking them. We see similar trends when using LLaMA-3.
To further confirm our findings, we conduct a qualitative study in the data-scarce setup, where we compute the ROUGE scores of the final 5 pseudo-labels for each method against the respective ground truth summaries, and Figure 4 shows the mean ROUGE-2 of the five selected pseudo-labels. To clarify, we obtained the “Oracle" results by directly selecting the final 5 pseudo-labels using ROUGE-2 scores computed against the ground truth. We note a stark difference between “Confidence" and “Oracle," which shows that relying solely on teacher confidence consistently leads to a selection of low-quality pseudo-labels. Combining GPT-4 score with teacher confidence is effective (“Confidence + G-4 score"), and adding the GPT-4 relabeling greatly boosts the quality of selected pseudo-labels (“Ours").
@@ 5.2 Effect of Relabeling
Referring to Tables 2 and 3, we observe that relabeling with LLMs leads to a significant boost in the summarization performance in terms of both ROUGE scores and L-Eval. When using BERTbase as the backbone, we note that the ROUGE-1 improves from 52.3 to 58.9 on the TweetSumm dataset, 25.2 to 26.1 on the WikiHow dataset, and 27.7 to 29.1 on the ArXiv/PubMed dataset. GPT-4 relabeling is also effective when using teacher confidence without GPT-4 score (“Confidence" v/s “Confidence + G-4 relabeling"). Our previous qualitative study also supports these results, showing that relabeling improves the quality of pseudo-labels. We observe similar trends when using DistilBERTbase as PreSumm’s backbone and LLaMA-3 instead of GPT-4. When using 500 labels, we note boosts in performance but the relative scale is smaller compared to when using 50 labels.
We conduct additional testing to analyze the performance of our best- and second-best-performing models, both of which involve relabeling. We find that the p-value < 0.016 for Welch’s test for an R-1 of 58.9 (1.4) for “Confidence + G-4 score + G-4 relabelling" v/s 57.8 (1.2) for “Confidence + G-4 relabeling" on the TweetSumm dataset, denoting the differences are significant.
@@ 5.3 On Sample Efficiency of Different Approaches
Referring to Table 3, we now analyze the sample efficiency of different learning algorithms. For fully supervised methods, we note that including more labeled examples improves L-Eval and ROUGE scores across the board (“PreSumm (50 labels)" v/s “PreSumm (300 labels)"). Our semi-supervised approach using 50 labels with GPT-4 relabeling and GPT-4 score achieves competitive performance to the fully supervised PreSumm model trained on 300 labels. Notably, we get better L-Eval scores than “PreSumm (300 labels)" on WikiHow and ArXiv/PubMed datasets and are competitive on TweetSumm. This is encouraging, as “PreSumm (300 labels)" approximates the best-case scenario of 100% high-quality labels in the training set. In the data-abundant setting, our proposed method with LLaMA outperforms the respective fully supervised model in terms of both ROUGE and L-Eval.
We plot the R-1 scores against the number of training cycles, as shown in Figure 5. Overall, “Random" setting is highly unstable, “Confidence + G-4 score" slightly improves over “Confidence" on TweetSumm and WikiHow, but more importantly, it is consistently more stable. Finally, our method with GPT-4 scoring and relabeling not only significantly boosts the R-1 scores (visible gap between “Ours" and the rest) but also does so at a much faster rate. For all the datasets, our method peaks and stabilizes under 20 cycles (100 pseudo-labels), further endorsing the sample efficiency of our method compared to other approaches.
@@ 6 Conclusion
First, we rely completely on LLMs pseudo-labels for scoring and relabelling, which is not perfect and we do not currently have a way to automatically judge them. We need a human in the loop to judge the quality of such pseudo-labels.
Second, we can also potentially improve the sample efficiency of the model. Currently, we relabel all the top 50 pseudo-labels selected by PreSumm in Stage 3. This is cheaper than pseudo-labeling all the pseudo-labels, but there’s a chance that LLM relabels an already good-quality summary, as we note from our analysis that the initial shortlist contains some good pseudo-labels. Therefore, we can develop an additional strategy to decide whether to relabel a given pseudo-label. Such a method would further reduce costs for using closed-source models like GPT-4. We hope this work furthers research toward incorporating LLMs to address such problems for low-resource semi-supervised tasks.
@@
We use GPT and LLaMA-3 to generate new examples, and even though they is instruction-tuned, some generations might depict undesirable biases with respect to the current objective. For the stated reason, we recommend using our proposed models with human monitoring when employing the method for real-world settings. Practitioners may also consider explicitly debiasing language models for their specific use cases (Barikeri et al., 2021; Schick et al., 2021).